Circular No. 9080

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2009 T1 (McNAUGHT) R. H. McNaught reports his discovery of a comet with a diffuse 14′′

coma and a diffuse 0′.6 tail in p.a. 145^{o} on CCD images taken with the 0.5-m Uppsala Schmidt telescope (discovery observation tabulated below). After posting on the Minor Planet Center’s ‘NEOCP’ webpage, other CCD astrometrists also have commented on the object’s cometary appearance. A. C. Gilmore and P. M. Kilmartin (Mt. John 1.0-m reflector, Oct. 10.7 UT) note a 30′′ tail in p.a. 140^{o}. R. Ligustri, Udine, Italy, reports that his co-added images (taken remotely with an RAS Observatory 0.40-m reflector near Moorook, Australia; Oct. 10.7) show a coma of diameter ∼ 12′′ and a thin 20′′ tail in p.a. 167^{o}. D. Chestnov, Moscow, writes that his images (taken remotely with a 0.15-m refractor at the Tzec Maun Observatory, near Pingelly, W. Australia; Oct. 10.7) show a small 10′′ coma with a short tail in p.a. 185^{o} and a nuclear condensation elongated in p.a. 140^{o}. C. Jacques and E. Pimentel (Belo Horizonte, MG, Brazil, 0.30-m reflector, Oct. 11.3) find an 18′′ coma in fifty 30-s stacked exposures.

2009 UT α_{2000} δ_{2000} Mag. Observer
Oct. 9.74900 6^{h}11^{m}14.^{s}33 −58^{o}50′27′′.8 17.8 McNaught

The available astrometry, the following preliminary parabolic orbital elements, and an ephemeris appear on MPEC 2009-T36.

T = 2010 Mar. 25.069 TT  ω = 300.^{o}574
Ω = 59.778 }2000.0
q = 5.56975 AU  i = 88.081

COMETS C/2009 G7, C/2009 G8, C/2009 H3 (SOHO) Further to IAUC 9079, additional presumed comet have been found

on SOHO website images. K. Battams reports that C/2009 G7, a Kreutz sungrazer, was stellar in appearance and tiny (mag ∼ 7) in C3 images, and small and condensed in C2 images. C/2009 G8, a non-group comet, was slightly diffuse and elongated (mag ∼ 7–7.5). C/2009 H3, also a Kreutz sungrazer, was diffuse and elongated (mag ∼ 8).

Comet 2009 UT α_{2000} δ_{2000} Inst. F MPEC
C/2009 G7 Apr. 13.488 1^{h}44.^{m}8 + 6^{o}34′ C3/2 AK 2009-N33
C/2009 G8  13.963 1 33.2 +11 20 C2  BZ 2009-N33
C/2009 H3  18.504 1 51.9 + 9 46 C2  MK 2009-N33
2009 October 11  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2009 CBAT  Daniel W. E. Green